OpenStudy (anonymous):

Find the following product: (x - 3)(x + 3)

OpenStudy (kirbykirby):

x^2-9

OpenStudy (anonymous):

wat about -6

OpenStudy (kirbykirby):

There's an easy property for this so you don't have to do distribution. If you have (a^2-b^2), then this = (a+b)(a-b) So.. since you have (x-3)(x+3), you find (x)^2 = x^2 and (3)^2=9... so x^2-9

OpenStudy (anonymous):

ok can you help wit another one

OpenStudy (kirbykirby):

Ok. sure